# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v4

@@ 1 Introduction and main results
${\rm div}(a({\boldsymbol{y}})\nabla u({\boldsymbol{y}}))\ =\ f\quad\text{in}\quad D,\quad u({\boldsymbol{y}})|_{\partial D}\ =\ 0,$ (1.2)
where $a({\boldsymbol{y}})$ is of the log-normal form
$a({\boldsymbol{y}})=\exp(b({\boldsymbol{y}}))\ \ {\text{with }}\ b({\boldsymbol{y}})=\sum_{j=1}^{\infty}y_{j}\psi_{j},$ (1.3)
$y_{j}$ are i.i.d. standard Gaussian random variables and $\psi_{j}\in L_{\infty}(D)$.
@@
A natural question arising is whether the known convergence rates are optimal and whether one can improve them. The aim of the present paper is to improve these known bounds in the case of higher spatial regularity of the solution of the parametric equation (1.2) with log-normal inputs (1.3). More precisely, we are interested in the problem of fully discrete simultaneous spatial-parametric linear collocation approximation of solutions $u$ to parametric PDEs (1.2) on bounded polygonal domain $D$ with log-normal random inputs (1.3) and its convergence rate, based on a finite number of its particular spatial-parametric values $u({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,u({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})$. The approximation error is measured by the norm of the Bochner space $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$ associated with the energy space $V$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$ (see Subsection 2.2 for definition of Bochner space). The present paper can be considered as a continuation of the works [7, 20, 25]. In [20, 25], we have investigated the problem of fully discrete multi-level collocation approximation of stochastic and parametric elliptic PDEs with log-normal inputs by using sparse-grid GPC Lagrange interpolation at the zeros of Hermite polynomials. In the recent paper [7], we have investigated semi-discrete collocation approximation of these equations by employing extended least squares sampling algorithms in Bochner spaces. In the present paper, we develop and combine the different techniques used in these papers for solving the formulated problem.
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\ \ \text{and}\ \ b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\},$
and satisfy the condition ${\boldsymbol{b}}_{i}\in\ell^{p_{i}}({\mathbb{N}})$, respectively, with $0<p_{1}\leq p_{2}<1$ and $p_{1}<2/3$. Let the numbers $\alpha$ and $\beta$ be defined by
@@
For arbitrary positive number $\varepsilon$, there exist positive constants $C$ and $C_{\varepsilon}$ such that for all $n\geq 2$, there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\mathbb{R}}^{\infty}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\begin{split}\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq 1/p_{2}-3/2,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ 1/p_{2}-3/2<\alpha<1/p_{2}-1,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/p_{2}-1,\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/p_{2}-1.\end{cases}\end{split}$ (1.5)
[…]
It is worth emphasizing that $S_{n}$ is a fully discrete multi-level collocation approximation method. At each level in $S_{n}$, the spatial component is based on finite element Lagrange interpolation associated with triangulations of the spatial domain $D$, while the parametric component is based on Hermite-Lagrange GPC interpolation in the case $\alpha\leq 1/p_{2}-3/2$, and on extended least squares sampling algorithms in the cases $\alpha>1/p_{2}-3/2$.
Following the approach in [7, 20, 25], the convergence rate results in (1.5) and constructions of the linear sampling algorithms $S_{n}$ realizing them, are obtained as consequences of applications to infinite-dimensional holomorphic functions of general results on multi-level linear sampling recovery in abstract Bochner spaces. In the case of small spatial regularity $\alpha\leq 1/p_{2}-3/2$, we use a modification of methods of sparse-grid Hermite-Lagrange GPC interpolation in Bochner spaces in [20]. In the cases of higher spatial regularity $\alpha>1/p_{2}-3/2$, we used extended least squares sampling algorithms in Bochner spaces and some recent results on linear sampling recovery of functions in reproducing kernel Hilbert spaces [8, 24, 28, 29]. We would like to emphasize that the general theory of sampling recovery in abstract Bochner spaces presented in this paper as well as [7, 20, 25] is applicable to uncertainty quantification for a wide class of parametric and stochastic PDEs.
@@
In Section 2, we investigate multi-level linear sampling recovery of functions in an abstract Bochner space $L_{2}(U,X;\mu)$ for a Hilbert space $X$ and a probability measure space $(U,\Sigma,\mu)$, based on some weighted $\ell_{2}$-summability of their coefficients of an orthonormal expansion, and some approximation properties in the space $X$. The received results are then applied to $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$, the Bochner space associated with a Hilbert space $X$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$. In Section 3, we apply the results on multi-level linear sampling recovery in the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ in the previous section to $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions on $\mathbb{R}$$^{\infty}$. In Section 4, we construct fully discrete multi-level sparse-grid sampling algorithms for $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions, based on GPC Lagrange-Hermite interpolation, and prove convergence rates of the approximation by them. In Section 5, we prove that the parametric solution $u$ to the parametric elliptic PDEs (1.1) on bounded polygonal domain with log-normal inputs (1.3) is a $({\boldsymbol{b}}_{j},\xi,\delta,V)$-holomorphic function. This allows prove convergence rates of a fully discrete multi-level collocation algorithm by applying the results in Sections 3 and 4. In Section 6, we discuss various least squares sampling algorithms for functions in the reproducing kernel Hilbert space, and inequalities between sampling $n$-widths and Kolmogorov $n$-widths of the unit ball of this space, and how to apply these inequalities to obtain corresponding convergence rates of multi-level linear sampling recovery in abstract Bochner spaces and of fully discrete multi-level collocation approximation of the parametric solution $u$ to the parametric elliptic PDEs (1.1) on bounded polygonal domain with log-normal inputs (1.3).
@@
Notation As usual, $\mathbb{N}$} denotes the natural numbers, $\mathbb{Z}$} the integers, $\mathbb{R}$} the real numbers, $\mathbb{C}$} the complex numbers, and ${\mathbb{N}}_{0}:=\{s\in{\mathbb{Z}}:s\geq 0\}$. We denote $\mathbb{R}$$^{\infty}$ the set of all sequences ${\boldsymbol{y}}=(y_{j})_{j\in{\mathbb{N}}}$ with $y_{j}\in{\mathbb{R}}$. Denote by $\mathbb{F}$} the set of all sequences of non-negative integers ${\boldsymbol{s}}=(s_{j})_{j\in{\mathbb{N}}}$ such that their support $\operatorname{supp}({\boldsymbol{s}}):=\{j\in{\mathbb{N}}:s_{j}>0\}$ is a finite set. If ${\boldsymbol{a}}=(a_{j})_{j\in{\mathcal{J}}}$ is a set of positive numbers with any index set $\mathcal{J}$}, then we use the notation ${\boldsymbol{a}}^{-1}:=(a_{j}^{-1})_{j\in{\mathcal{J}}}$. We use letters $C$ and $K$ to denote general positive constants which may take different values, and $C_{a,b,...,d}$ and $K_{a,b,...,d}$ constants depending on a,b,,da,b,...,d. For the quantities $A_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})$ depending on $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$, we write $A_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ ($n\in{\mathbb{N}}$ is specially dropped), if there exists some constant $C>0$ such that $A_{n}(f,{\boldsymbol{k}})\leq CB_{n}(f,{\boldsymbol{k}})$ for all $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ (the notation $A_{n}(f,{\boldsymbol{k}})\gg B_{n}(f,{\boldsymbol{k}})$ has the obvious opposite meaning), and $A_{n}(f,{\boldsymbol{k}})\asymp B_{n}(f,{\boldsymbol{k}})$ if $A_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})\ll A_{n}(f,{\boldsymbol{k}})$. Denote by $|G|$ the cardinality of the set $G$.
@@ 2.1 Function spaces
Let $(U,{\mathcal{A}},\mu)$ be a probability measure space with $U$ being a separable topological space and let $X$ be either a separable complex Hilbert space or $\mathbb{C}$}. Denote by $L_{2}(U,X;\mu)$ the Bochner space of strongly $\mu$-measurable mappings $v$ from $U$ to $X$, equipped with the norm
@@
The space $L_{2}(U,{\mathbb{C}};\mu)$ consists of all square $\mu$-integrable complex functions on $U$. Since $U$ separable, the spaces $L_{2}(U,{\mathbb{C}};\mu)$ and $L_{2}(U,X;\mu)$ are also separable complex Hilbert spaces, and moreover,
@@
Let $\left(\varphi_{s}\right)_{s\in{\mathbb{N}}}$ be a fixed orthonormal basis of $L_{2}(U,{\mathbb{C}};\mu)$. Then a function $v\in L_{2}(U,X;\mu)$ can be represented by the expansion
@@
In particular, the space $H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ is the subspace in $L_{2}(U,{\mathbb{C}};\mu)$ equipped with its own inner product
@@
[…]
@@
Let $X\not$$\mathbb{C}$} and $(V_{k})_{k\in{\mathbb{N}}_{0}}$ be a given sequence of subspaces $V_{k}\subset X$ of dimension $2^{k}$, and $(P_{k}^{X})_{k\in{\mathbb{N}}_{0}}$ a given sequence of uniformly bounded linear projectors $P_{k}^{X}$ from $X$ onto $V_{k}$. We extend the operators $P_{k}^{X}$ as uniformly bounded linear operators in $L_{2}(U,X;\mu)$ (which with an abuse is denoted again by $P_{k}^{X}$) such that
@@
For $i=1,2$, let $0<q_{i}\leq 2$ and ${\boldsymbol{\sigma}}_{i}:=(\sigma_{i;s})_{s\in{\mathbb{N}}}$ be given non-decreasing sequences of numbers strictly larger than $1$, such that ${\boldsymbol{\sigma}}_{i}^{-1}:=\left(\sigma_{i,s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{q_{i}}({\mathbb{N}})$. For a given number $\alpha>0$, denote by $H_{X}^{\alpha}$ the linear subspace in $L_{2}(U,X;\mu)$ of all $v$ such that the norm
@@
$f\mapsto\sum_{k\in{\mathbb{N}}}\left(\sum_{s\in{\mathbb{N}}}a_{k,s}f_{k}\right)\varphi_{k}.$ (2.9)
@@
These correspondence and norm equality between bounded linear operators $A^{X}$ and $A^{\mathbb{C}}$ allows us to define the spaces $H_{\mathbb{C}}^{\alpha}$ and $H_{{\mathbb{C}}}^{\alpha,\tau}$ generated from given spaces $H_{X}^{\alpha}$ and $H_{X}^{\alpha,\tau}$, by replacing $P_{k}^{X}$ with $P_{k}^{\mathbb{C}}$ and $H_{X,{\boldsymbol{\sigma}}_{i}}$ with $H_{{\mathbb{C}},{\boldsymbol{\sigma}}_{i}}$, $i=1,2$, in their definitions (2.4)–(2.5) and (2.6)–(2.7), respectively. For simplicity, in what follows, we will use the abbreviations $P_{k}:=P_{k}^{X}$ and $P_{k}:=P_{k}^{\mathbb{C}}$, provided this does not lead to any misunderstanding.
@@
$\|v-P_{k}v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}:=\ \left(\sum_{s\in{\mathbb{N}}}\left(\sigma_{s}\|(v-P_{k}v)_{s}\|_{X}\right)^{2}\right)^{1/2}<\infty.$
@@
$\begin{split}\sum_{s\in{\mathbb{N}}}\|v_{s}\varphi_{s}\|_{L_{2}(U,X;\mu)}\ &=\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\|\varphi_{s}\|_{L_{2}(U,{\mathbb{C}};\mu)}\ =\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\\ &\leq\bigg(\sum_{s\in{\mathbb{N}}}(\sigma_{s}\|v_{s}\|_{X})^{2}\bigg)^{1/2}\bigg(\sum_{s\in{\mathbb{N}}}\sigma_{s}^{-2}\bigg)^{1/2}\ <\ \infty.\end{split}$

### v5

@@ 1 Introduction and main results
${\rm div}(a({\boldsymbol{y}})\nabla u({\boldsymbol{y}}))\ =\ f\quad\text{in}\quad D,\quad u({\boldsymbol{y}})|_{\partial D}\ =\ 0.$ (1.2)
The present paper considers the case of parametric equation (1.2) when the random diffusion coefficient $a$ is of the log-normal form
$a({\boldsymbol{y}})=\exp(b({\boldsymbol{y}})),\quad{\text{with }}\ b({\boldsymbol{y}})=\sum_{j=1}^{\infty}y_{j}\psi_{j},$ (1.3)
where $y_{j}$ are i.i.d. standard Gaussian random variables and where $\psi_{j}\in L_{\infty}(D)$.
@@
A natural question arising is whether the known convergence rates are optimal and whether one can improve them. The aim of the present paper is to improve these known bounds in the case of higher spatial regularity of the solution of the parametric equation (1.2) with log-normal inputs (1.3). More precisely, we are interested in the problem of fully discrete simultaneous spatial-parametric linear collocation approximation of solutions $u$ to parametric PDEs (1.2) on bounded polygonal domain $D$ with log-normal random inputs (1.3) and its convergence rate, based on a finite number of its particular spatial-parametric values $u({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,u({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})$. The approximation error is measured by the norm of the Bochner space $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$ associated with the energy space $V$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$ (see Subsection 2.2 for definition of Bochner space). The present paper can be considered as a continuation of the works [7, 20, 25]. In [20, 25], we have investigated the problem of fully discrete multi-level collocation approximation of stochastic and parametric elliptic PDEs with log-normal inputs by using sparse-grid GPC Lagrange interpolation at the zeros of Hermite polynomials. In the recent paper [7], we have investigated semi-discrete approximation of these equations by employing extended least squares sampling algorithms in Bochner spaces. In the present paper, we develop and combine the different techniques used in these papers for solving the formulated problem.
@@
$b_{1;j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\ \ \text{and}\ \ b_{2;j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\},$
and satisfy the condition ${\boldsymbol{b}}_{i}\in\ell^{p_{i}}({\mathbb{N}})$, respectively, with $0<p_{1}\leq p_{2}<1$. Let the numbers $\alpha$ and $\beta$ be defined by
@@
For any fixed positive number $\varepsilon$, there exist positive constants $C$ and $C_{\varepsilon}$ such that for all $n\geq 2$, there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\mathbb{R}}^{\infty}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\begin{split}\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq{\color[rgb]{0,0,0}{1/p_{2}-1}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ {\color[rgb]{0,0,0}{1/p_{2}-1<\alpha<1/p_{2}-1/2}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha={\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\quad&{\rm if}\ \alpha>{\color[rgb]{0,0,0}{1/p_{2}-1/2}}.\end{cases}\end{split}$ (1.5)
The convergence rates in (1.5) significantly improve the best-known convergence rates [25, Theorem 7.5 and page 177] of fully discrete multilevel Smolyak sparse-grid interpolation. Moreover, in contrast to the previous works, in the particular intermediate case $1/p_{2}-3/2<\alpha\leq 1/p_{2}-1/2$, the maximal convergence rate $n^{-\alpha}$ still holds with a logarithm factor. It is important to highlight that with some logarithm factors, the convergence rates in (1.5) coincide with the convergence rates $n^{-\min(\alpha,\beta)}$ of fully discrete best $n$-term approximation of the parametric solution $u$ in the space $L_{2}({\mathbb{R}}^{\infty},V;\gamma)$. See Remark 5.5 for a detailed commentary on the related results in prior works and contrast of them with the main findings (1.5) of this paper.
It is worth emphasizing that $S_{n}$ is a fully discrete multi-level collocation approximation method. At each level in $S_{n}$, the spatial component is based on finite element Lagrange interpolation associated with triangulations of the spatial domain $D$, while the parametric component is based on Hermite-Lagrange GPC interpolation in the case $\alpha\leq 1/p_{2}-1$ , and on extended least squares sampling algorithms in the cases $\alpha>1/p_{2}-1$ .
Following the approach in [7, 20, 25], the convergence rate results in (1.5) and constructions of the linear sampling algorithms $S_{n}$ realizing them, are obtained as consequences of applications to infinite-dimensional holomorphic functions of general results on multi-level linear sampling recovery in abstract Bochner spaces. In the case of small spatial regularity $\alpha\leq 1/p_{2}-1$ , we use a modification of methods of sparse-grid Hermite-Lagrange GPC interpolation in Bochner spaces in [20]. In the cases of higher spatial regularity $\alpha>1/p_{2}-1$ , we used extended least squares sampling algorithms in Bochner spaces and some recent results on linear sampling recovery of functions in reproducing kernel Hilbert spaces [8, 24, 28, 29]. We would like to emphasize that the general theory of sampling recovery in abstract Bochner spaces presented in this paper as well as [7, 20, 25] is applicable to uncertainty quantification for a wide class of parametric and stochastic PDEs.
@@
In Section 2, we investigate multi-level linear sampling recovery of functions in an abstract Bochner space $L_{2}(U,X;\mu)$ for a Hilbert space $X$ and a probability measure space $(U,{\mathcal{A}},\mu)$ , based on some weighted $\ell_{2}$-summability of their coefficients of an orthonormal expansion, and some approximation properties in the space $X$. The received results are then applied to $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$, the Bochner space associated with a Hilbert space $X$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$. In Section 3, we apply the results on multi-level linear sampling recovery in the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ in the previous section to $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions on $\mathbb{R}$$^{\infty}$. In Section 4, we construct fully discrete multi-level sparse-grid sampling algorithms for $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions, based on GPC Lagrange-Hermite interpolation, and prove convergence rates of the approximation by them. In Section 5, we prove that the parametric solution $u$ to the parametric elliptic PDEs (1.1) on bounded polygonal domain with log-normal inputs (1.3) is a $({\boldsymbol{b}}_{j},\xi,\delta,V)$-holomorphic function. This allows to prove convergence rates of a fully discrete multi-level collocation algorithm by applying the results in Sections 3 and 4. In Section 6, we discuss various least squares sampling algorithms for functions in the reproducing kernel Hilbert space, and inequalities between sampling $n$-widths and Kolmogorov $n$-widths of the unit ball of this space, and how to apply these inequalities to obtain corresponding convergence rates of multi-level linear sampling recovery in abstract Bochner spaces and of fully discrete multi-level collocation approximation of the parametric solution $u$ to the parametric elliptic PDEs (1.1) on bounded polygonal domain with log-normal inputs (1.3).
@@
Notation As usual, $\mathbb{N}$} denotes the natural numbers, $\mathbb{Z}$} the integers, $\mathbb{R}$} the real numbers, $\mathbb{C}$} the complex numbers, and ${\mathbb{N}}_{0}:=\{s\in{\mathbb{Z}}:s\geq 0\}$. We denote $\mathbb{R}$$^{\infty}$ the set of all sequences ${\boldsymbol{y}}=(y_{j})_{j\in{\mathbb{N}}}$ with $y_{j}\in{\mathbb{R}}$. Denote by $\mathbb{F}$} the set of all sequences of non-negative integers ${\boldsymbol{s}}=(s_{j})_{j\in{\mathbb{N}}}$ such that their support $\operatorname{supp}({\boldsymbol{s}}):=\{j\in{\mathbb{N}}:s_{j}>0\}$ is a finite set. If ${\boldsymbol{a}}=(a_{j})_{j\in{\mathcal{J}}}$ is a set of positive numbers with any index set $\mathcal{J}$}, then we use the notation ${\boldsymbol{a}}^{-1}:=(a_{j}^{-1})_{j\in{\mathcal{J}}}$. We use letters $C$ and $K$ to denote general positive constants which may take different values, and $C_{a,b,...,d}$ and $K_{a,b,...,d}$ constants depending on a,b,,da,b,...,d. Denote by $|G|$ the cardinality of the set $G$.
@@ 2.1 Function spaces
Let $(U,{\mathcal{A}},\mu)$ be a probability measure space with $U$ being a separable topological space and let $X$ be a separable complex Hilbert space. Denote by $L_{2}(U,X;\mu)$ the Bochner space of strongly $\mu$-measurable mappings $v$ from $U$ to $X$, equipped with the norm
@@
Denote by $L_{2}(U,{\mathbb{C}};\mu)$ the Hilbert space which consists of all square $\mu$-integrable complex functions on $U$. Since $U$ is separable, the spaces $L_{2}(U,{\mathbb{C}};\mu)$ and $L_{2}(U,X;\mu)$ are also separable complex Hilbert spaces, and moreover,
@@
Let $\left(\varphi_{s}\right)_{s\in{\mathbb{N}}}$ be a fixed orthonormal basis of $L_{2}(U,{\mathbb{C}};\mu)$. Then every function $v\in L_{2}(U,X;\mu)$ can be represented by the expansion
@@
Similarly, the space $H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ is the subspace in $L_{2}(U,{\mathbb{C}};\mu)$ equipped with its own inner product
@@
having the eigenfunctions $\left(\varphi_{s}\right)_{s\in{\mathbb{N}}}$ and the eigenvalues $\left(\sigma_{s}\right)_{s\in{\mathbb{N}}}$ . Moreover, $K({\boldsymbol{x}},{\boldsymbol{y}})$ satisfies the finite trace assumption
@@
Let $(V_{k})_{k\in{\mathbb{N}}_{0}}$ be a given sequence of subspaces $V_{k}\subset X$ of dimension $2^{k}$, and $(P_{k}^{X})_{k\in{\mathbb{N}}_{0}}$ a given sequence of uniformly bounded linear projectors $P_{k}^{X}$ from $X$ onto $V_{k}$. We extend the operators $P_{k}^{X}$ as uniformly bounded linear operators in $L_{2}(U,X;\mu)$ (which with an abuse is denoted again by $P_{k}^{X}$) such that
@@
For $i=1,2$, let $0<q_{i}\leq 2$ and ${\boldsymbol{\sigma}}_{i}:=(\sigma_{i;s})_{s\in{\mathbb{N}}}$ be given non-decreasing sequences of numbers strictly larger than $1$, such that ${\boldsymbol{\sigma}}_{i}^{-1}:=\left(\sigma_{i;s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{q_{i}}({\mathbb{N}})$. For a given number $\alpha>0$, denote by $H_{X}^{\alpha}$ the linear subspace in $L_{2}(U,X;\mu)$ of all $v$ such that the norm
@@
Note that in the definition (2.7), the subspaces $(V_{k})_{k\in{\mathbb{N}}_{0}}$ is not required to be nested and the projection operators $(P_{k}^{X})_{k\in{\mathbb{N}}_{0}}$ not to be required orthogonal.
@@
$f\mapsto\sum_{k\in{\mathbb{N}}}\left(\sum_{s\in{\mathbb{N}}}a_{k,s}{\color[rgb]{0,0,0}{f_{s}}}\right)\varphi_{k}.$ (2.9)
@@
These correspondence and norm equality between bounded linear operators $A^{X}$ and $A^{\mathbb{C}}$ allows us to define the spaces $H_{\mathbb{C}}^{\alpha}$ and $H_{{\mathbb{C}}}^{\alpha,\tau}$ generated from given spaces $H_{X}^{\alpha}$ and $H_{X}^{\alpha,\tau}$, by replacing $P_{k}^{X}$ with $P_{k}^{\mathbb{C}}$ and $H_{X,{\boldsymbol{\sigma}}_{i}}$ with $H_{{\mathbb{C}},{\boldsymbol{\sigma}}_{i}}$, $i=1,2$, in their definitions (2.4)–(2.5) and (2.6)–(2.7), respectively.
For convenience, in what follows, we will use the letter $X$ to denote a separable Hilbert space or the complex plane $\mathbb{C}$}, and use the abbreviations $P_{k}:=P_{k}^{X}$ and $P_{k}:=P_{k}^{\mathbb{C}}$, provided this does not lead to any misunderstanding.
@@
$\|v-P_{k}v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}:=\ \left(\sum_{s\in{\mathbb{N}}}\left({\color[rgb]{0,0,0}{\sigma_{2;s}}}\|(v-P_{k}v)_{s}\|_{X}\right)^{2}\right)^{1/2}<\infty.$
@@
$\begin{split}\sum_{s\in{\mathbb{N}}}\|v_{s}\varphi_{s}\|_{L_{2}(U,X;\mu)}\ &=\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\|\varphi_{s}\|_{L_{2}(U,{\mathbb{C}};\mu)}\ =\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\\ &\leq\bigg(\sum_{s\in{\mathbb{N}}}({\color[rgb]{0,0,0}{\sigma_{2;s}}}\|v_{s}\|_{X})^{2}\bigg)^{1/2}\bigg(\sum_{s\in{\mathbb{N}}}{\color[rgb]{0,0,0}{\sigma_{2;s}}}^{-2}\bigg)^{1/2}\ <\ \infty.\end{split}$